OpenStudy (anonymous):

A student collected the following data for a fixed volume of gas: Temperature (⁰C) Pressure (mm of Hg) 10 726 20 750 40 800 70 880 100 960 150 ??? Fill in the missing data point. Show all calculations leading to an answer.

OpenStudy (rushwr):

we know for a fixed volume the no. of moles is also fixed. SO in PV=nRT equation V , n and R is constant ( R is universal gas constant so it is anyways constant ) So we can say that \[P \alpha T\] Pressure is directly proportional to the temperature. \[\frac{ P _{1} }{ T _{1} }= \frac{ P _{2} }{ T _{2} }\]

OpenStudy (anonymous):

okay can please explain more

OpenStudy (rushwr):

So use one set of known of data and the unkonwn set like this ! P1/T1=P2/T2 So let's say that the unknown pressure is P1 then P1= P2 x T1 /T2

OpenStudy (rushwr):

If we use the 1st set of data given ! P1= ? T1= 150 C P2 = 726 T2 = 10

OpenStudy (rushwr):

Did u get it ?

OpenStudy (anonymous):

yes thanks